# VITA: Vision-to-Action Flow Matching Policy

### v2

@@ 4.2 VITA Architecture Design
Figure 2: An overview of the VITA architecture: The vision encoder maps observations into a source latent representation $\bm{z}_{0}$ for the flow; the action encoder provides a target latent representation $\bm{z}_{1}$ for flow matching training. The action decoder learns to decode $\hat{\bm{z}}_{1}$ (latent actions generated by solving ODEs) to actions via flow latent decoding losses, and decode $\bm{z}_{1}$ to actions (latent actions from action encoder) via AE losses. The flow matching network is designed to learn the velocity field over a continuous flow matching path from $\bm{z}_{0}$ to $\bm{\hat{z}}_{1}$.
As depicted in Figure˜2, VITA  is composed of the three primary components: 1) The Vision Encoder ($\mathcal{E}_{v}$) maps raw camera images into a vector representation of visual latents $I$. The observation $O$ to the policy consists of both image vector representations $I$ and, optionally, the robot’s proprioceptive states $S$. In short, $O$ is mapped into a latent vector $\bm{z}_{0}=\mathcal{E}_{v}(O),\quad\text{where }\bm{z}_{0}\in\mathbb{R}^{D_{\text{latent}}}$ is the source of flow matching. 2) The Action Autoencoder (AE) consists of the Action Encoder and the Action Decoder, and learns a compact representation for action chunks. The Action Encoder ($\mathcal{E}_{a}$) maps the ground-truth action chunk $A$ to latent actions $\bm{z}_{1}=\mathcal{E}_{a}(A),\quad\text{where }\bm{z}_{1}\in\mathbb{R}^{D_{\text{latent}}}$ serves as the target for flow matching; the Action Decoder ($\mathcal{D}_{a}$) reconstructs an action chunk $\hat{A}=\mathcal{D}_{a}(\hat{\bm{z}}_{1})$ from latent actions $\hat{\bm{z}}_{1}$. 3) The Flow Matching Network ($v_{\theta}$) is learned to predict the velocity field at arbitrary $t$.
@@ 4.3 Bridging the Modality Gap between Vision and Action
A key constraint of flow matching is that the source and target distributions must share the same dimensionality. This poses a critical challenge for vision-to-action policies, since action spaces are typically much lower-dimensional than visual representations. For example, action dimensionalities range from 2 on PushT to 21 on ThreadNeedle, whereas visual representations can be 512-dimensional (zhao2024aloha, ) or even higher when using grid-based tokenizers (chi2023diffusion, ).
To bridge the gap, one naive option is to down-sample latent visual representations to action chunk dimensionalities which, however, causes severe information loss and degrades performance. Alternatively, one can up-sample actions with zero-padding, yielding sparse, unstructured targets that hinder flow matching learning (see Appendix B.1). A third alternative is a pre-trained, frozen action AE, akin to common practice in latent diffusion for image generation (stablediffusion, ), but this proves ineffective for learning vision-to-action flow: with sparse, limited action data the induced latent space is unreliable as a flow target and cannot be corrected once frozen. As another alternative, jointly training the action AE with flow matching may still fail, and our empirical studies identify the root cause as latent space collapse induced by a training-inference gap in the latent actions used for decoding as detailed below.
@@
During training, the decoder reconstructs actions from encoder-based latent actions $\bm{z}_{1}$, whereas at inference it decodes $\hat{\bm{z}}_{1}$ generated by solving the flow matching ODE. Since $\hat{\bm{z}}_{1}$ is an approximation, and does not always align with $\bm{z}_{1}$, the decoder can fail to map them into meaningful actions. To address this gap, we propose flow latent decoding, which enforces the model to decode from ODE-generated latent actions $\hat{\bm{z}}_{1}$ during training, anchoring the latent generation process with ground-truth actions.
@@ 4.4 VITA Learning Objectives
Under ˜1, for any $\hat{\bm{z}}_{1}$ in the neighborhood,
$m\,\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|-\varepsilon_{\mathrm{AE}}\ \leq\ \|\mathcal{D}_{a}(\hat{\bm{z}}_{1})-A\|\ \leq\ L\,\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|+\varepsilon_{\mathrm{AE}}.$
@@
We will provide further discussion in Section˜5.3, showing that including the flow latent decoding loss with a non-zero $\lambda_{\text{FLD}}$ is critical for avoiding latent space collapse and training successful VITA policies. We will also ablate the effects of $\lambda_{\text{FLD}}$, $\lambda_{\text{FLC}}$, and $\lambda_{\text{AE}}$ in Figure˜6.
@@
The training objective can be written as a weighted sum of all three losses: $\mathcal{L}_{\text{VITA}}=\lambda_{\text{FM}}\mathcal{L}_{\text{FM}}+\lambda_{\text{FLD}}\mathcal{L}_{\text{FLD}}+\lambda_{\text{AE}}\mathcal{L}_{\text{AE}}$.
@@ 5 Experiments
We evaluate VITA on ALOHA (chuang2024active, ; fu2024mobile, ) across 5 simulation tasks (Figure˜3) and 2 real-world tasks (Figure˜4). The simulation environments and datasets are from AV-ALOHA (chuang2024active, ), which extend ALOHA with an additional 7-DoF arm equipped with an active-vision camera that dynamically adjusts its viewpoint. This setup increases task difficulty due to the larger action space and non-stationary observation dynamics. In addition, we evaluate on 2 single-arm Robomimic simulation tasks.
@@
For simulation tasks, each environment provides 100-200 demonstrations. Demonstrations were collected from human experts in simulation using VR headsets, with the left-eye image used as input. VITA was trained at $25/3$ FPS on AV-ALOHA and $20$ FPS on Robomimic. For real-world tasks on AV-ALOHA, (Figure˜4). Each task is trained with 50 demonstrations, using the left image from the stereo active-vision camera as input. VITA was trained at 11 FPS, and linearly interpolated to 33 FPS for deployment on the physical robot.
Figure 3: An illustration of 5 AV-ALOHA tasks (CubeTransfer, SlotInsertion, HookPackage, PourTestTube, ThreadNeedle), and 2 Robomimic tasks (Square, Can).
@@
Figure 4: An illustration of two challenging real-world AV-ALOHA tasks, HiddenPick, and TransferFromBox. The pictures are taken from autonomous rollouts by the VITA policy.
5.1 Implementation
@@
Flow Matcher. We adopt OT-CFM (ot-cfm, ), a conditional flow matching method based on optimal transport, and solve the Euler ODE with 6 steps using linearly interpolated $t$ for both VITA and FM baselines.
Networks. We use the widely adopted ResNet-18 (resnet, ; su2025dense, ; chi2023diffusion, ) as the vision encoder for all the experiments. We use the 512-dimensional feature vector after the average pooling layer as the visual representations for VITA. Therefore, the source $\bm{z}_{0}$ and target latents $\bm{z}_{1}$ are both 512-dimensional vectors ($D_{\text{latent}}=512$). The vector representations and conditioning-free design is amenable to simple and efficient implementation. For example, we use 4-layer MLP-only networks for both flow matching and decoding.
Training & Evaluation. We train VITA on each task for 50,000 steps with a batch size of 128. The policy predicts action chunks of length 16, of which the first 8 actions are executed. In simulation, we evaluate the policy every 500 steps, with each evaluation consisting of 50 episodes. We report the best success rates (SRs), averaged over 3 random seeds, for both VITA and baselines (Table˜1). For real-world tasks, we evaluate 3 checkpoints over 20 episodes each and report the best SRs (Table˜2).
Baselines. We compare VITA against state-of-the-art generative policies, including flow matching (FM) policy (fmp-affordance, ), diffusion policy (DP)(chi2023diffusion, ), and action chunking transformer (ACT)(zhao2023learning, ). We follow the ACT and DP implementation in LeRobot cadene2024lerobot . We found DP requires significantly more training steps than FM; ACT yields sub-optimal performance on most tasks; therefore, we increase training steps to 200,000 for both DP and ACT, and then report the best success rates. For DP, each inference takes 10 DDPM (Denoising Diffusion Probabilistic Model) steps (chi2023diffusion, ). We evaluate both the task performance and inference efficiency of VITA relative to these baselines in Section˜5.2.
@@ 5.2 Performance
Success Rates.
We evaluate VITA against state-of-the-art conditional generative policies on 8 simulated tasks: 5 bimanual AV-ALOHA tasks with active vision (chuang2024active, ), 2 single-arm Robomimic tasks (robomimic2021, ), and 1 PushT task. Results are reported in Table˜1. Across all simulation tasks, VITA consistently outperforms or matches state-of-the-art baselines. ACT shows limited performance even after convergence, while DP requires significantly more training steps and remains sub-optimal despite extended training. We further assess VITA in real-world settings on two challenging AV-ALOHA tasks, with performance summarized in Table˜2 and examples of autonomous rollouts in Figure˜4, highlighting the precision of vision-to-action flow in manipulation tasks with high-dimensional visual inputs and large action spaces.
@@
Table 1: Task success rates (SR) across simulation tasks including ALOHA, Robomimic, and PushT. We report the best SR during validation, and compute the mean across 3 random seeds.
@@
ThreadNeedle 0.92 0.88 0.00 0.52
SlotInsertion 0.76 0.83 0.20 0.50
PourTestTube 0.79 0.84 0.06 0.14
HookPackage 0.86 0.83 0.12 0.22
CubeTransfer 1.00 1.00 0.76 0.98
PushT 0.88 0.84 0.80 0.30
Square 1.00 1.00 1.00 0.46
Can 1.00 1.00 1.00 0.78
Table 2: Success rates of VITA on two real-world AV-ALOHA tasks. Each task has 3 subtasks.
@@
Subtask Reveal Pick Place Pick Transfer Place
SR 1.00 0.65 0.65 1.00 0.95 0.90
Efficiency.
The VITA policy is highly efficient due to its architectural simplicity. The core components used during inference, the flow matching network and the action decoder, are both lightweight MLP-only networks. We use vector representations for both vision and action, further reducing time and space overhead. VITA contains 31M parameters, including the ResNet-18. We benchmark the inference latency and policy throughput of VITA and compare it against conventional flow matching policy baselines. We implement two different flow matching parameterizations, including diffusion transformer (dit, ), and U-Net (unet, ), and three different conditioning mechanisms, including AdaLN (dit, ; dit-policy, ), cross-attention (attention, ; dit-policy, ; carp, ), and FiLM (perez2018film, ; chi2023diffusion, ), as the baselines. FM and DP are typically parameterized using U-Nets (unet, ) or diffusion transformers (DiTs) (dit, ), and are trained to predict velocity fields or noise at each denoising step. While U-Nets and DiTs are highly expressive, they can be large or computationally expensive, posing a drawback for real-world robotic deployments that require low-latency inference. Most importantly, these methods inevitably incorporate visual conditioning mechanisms, such as cross, further increasing both inference time and memory footprint.
With a control frequency of $25/3\approx 8.33$ Hz, our MLP-only VITA policy achieves an inference wall-clock time of 0.22 ms on an NVIDIA 4090 GPU. In contrast, a comparable transformer-based flow matching policy incurs higher latency (0.33-0.51 ms) and requires more parameters. A detailed breakdown of efficiency metrics is given in Table˜3.
Table 3: Model sizes and inference latency of different policy implementations.
Model Conditioning Architecture Num of Params Latency (ms/chunk)
VITA N/A MLP 31.09M 0.2215
FM AdaLN DiT 31.16M 0.3307
FM Cross-Attn DiT 29.06M 0.5102
FM FiLM U-Net 84.05M 0.3650
DDPM FiLM U-Net 81.82M 2.5985

### v3

@@ 4.2 VITA Architecture Design
Figure 2: An overview of the VITA architecture: The vision encoder maps observations into a source latent representation $\bm{z}_{0}$ for the flow; the action encoder provides a target latent representation $\bm{z}_{1}$ for flow matching training. The action decoder learns to decode $\hat{\bm{z}}_{1}$ (latent actions generated by solving ODEs) to actions via flow latent decoding losses, and decode $\bm{z}_{1}$ to actions (latent actions from action encoder) via autoencoder losses. The flow matching network learns the velocity field over a continuous flow matching path from $\bm{z}_{0}$ to $\bm{z}_{1}$.
As depicted in Figure˜2, VITA  is composed of three primary components: 1) The Vision Encoder first ($\mathcal{E}_{v}$) maps raw camera images into the image representation $I$. The observation $O$ to the policy consists of $I$ and, optionally, the robot’s proprioceptive states $S$. Then, $O$ is mapped into a latent $\bm{z}_{0}=\mathcal{E}_{v}(O),\quad\text{where }\bm{z}_{0}\in\mathbb{R}^{D_{\text{latent}}}$ is the source of flow matching. 2) The Action Autoencoder (AE) consists of the Action Encoder and the Action Decoder, and learns a compact representation for action chunks. The Action Encoder ($\mathcal{E}_{a}$) maps the ground-truth action chunk $A$ to latent actions $\bm{z}_{1}=\mathcal{E}_{a}(A),\quad\text{where }\bm{z}_{1}\in\mathbb{R}^{D_{\text{latent}}}$ serves as the target for flow matching; the Action Decoder ($\mathcal{D}_{a}$) reconstructs an action chunk $\hat{A}=\mathcal{D}_{a}(\hat{\bm{z}}_{1})$ from latent actions $\hat{\bm{z}}_{1}$. 3) The Flow Matching Network ($v_{\theta}$) is learned to predict the velocity field at arbitrary $t$.
@@ 4.3 Bridging the Modality Gap between Vision and Action
A key constraint of flow matching is that the source and target distributions must share the same dimensionality. This poses a critical challenge for vision-to-action policies, since action spaces are typically much lower-dimensional than visual representations. For example, action dimensionalities range from 2 on PushT to 21 on ThreadNeedle, whereas visual representations can be 512-dimensional (zhao2024aloha) or even higher when using grid-based features.
To bridge the gap, one naive option is to down-sample latent visual representations to action chunk dimensionalities which, however, causes severe information loss and degrades performance. Alternatively, one can up-sample actions with zero-padding, yielding sparse, unstructured targets that hinder flow matching learning (see Appendix B.1). A third alternative is a pre-trained, frozen action AE, akin to common practice in latent diffusion for image generation (stablediffusion), but this proves ineffective for learning vision-to-action flow: with sparse, limited action data the induced latent space is unreliable as a flow target and cannot be corrected once frozen (see Appendix B.2). As another alternative, jointly training the action AE with flow matching may still fail, and our empirical studies identify the root cause as latent space collapse induced by a training-inference gap in the latent actions used for decoding as detailed below.
@@
During training, the decoder reconstructs actions from encoder-based latent actions $\bm{z}_{1}$, whereas at inference it decodes $\hat{\bm{z}}_{1}$ generated by solving the flow matching ODE. Since $\hat{\bm{z}}_{1}$ is an approximation and does not always align with $\bm{z}_{1}$, the decoder can fail to map them into meaningful actions. To address this gap, we propose flow latent decoding, which enforces the model to decode from ODE-generated latent actions $\hat{\bm{z}}_{1}$ during training, anchoring the latent generation process with ground-truth actions.
@@ 4.4 VITA Learning Objectives
Under Assumption˜1, for any $\hat{\bm{z}}_{1}$ in the neighborhood, we have $m\,\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|-\varepsilon_{\mathrm{AE}}\ \leq\ \|\mathcal{D}_{a}(\hat{\bm{z}}_{1})-A\|\ \leq\ L\,\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|+\varepsilon_{\mathrm{AE}}$.
@@
We further discuss this in Section˜5.3, showing that including the flow latent decoding loss with a non-zero $\lambda_{\text{FLD}}$ is critical for avoiding latent space collapse and training successful VITA policies. We will also ablate the effects of $\lambda_{\text{FLD}}$, $\lambda_{\text{FLC}}$, and $\lambda_{\text{AE}}$ in Figure˜6.
@@
The training objective is a weighted sum of all three losses: $\mathcal{L}_{\text{VITA}}=\lambda_{\text{FM}}\mathcal{L}_{\text{FM}}+\lambda_{\text{FLD}}\mathcal{L}_{\text{FLD}}+\lambda_{\text{AE}}\mathcal{L}_{\text{AE}}$.
@@ 5 Experiments
We evaluate VITA on 9 simulation and 5 real-world tasks, both spanning bimanual and single-arm manipulation. The bimanual tasks include 5 simulation (Figure˜3) and 2 real-world tasks (Figure˜4) on AV-ALOHA  (chuang2024active), which augments ALOHA  (zhao2024aloha) with an active-vision camera mounted on an additional 7-DoF arm. This challenging suite features high precision requirements, non-stationary observations using active vision, and 21-DoF high-dimensional actions. The single-arm tasks include 3 real-world tasks using one ALOHA arm (Figure˜4), featuring high randomization in object positions or colors, and 4 simulated tasks, including 2 Robomimic tasks (7D actions), PushT (2D actions), and CloseBox (9D actions) from RLBench (james2020rlbench).
@@
For simulation tasks, each environment provides 100-200 demonstrations. AV-ALOHA demonstrations were collected via expert teleoperation in VR, using the left-eye image as input; single-arm ALOHA demonstrations were collected using a leader arm. For the remaining tasks, we use publicly available datasets. In real-world AV-ALOHA experiments, each task is trained from 50 demonstrations using the left stereo image, and the single-arm ALOHA tasks are trained from 50-100 demonstrations using the wrist camera and an overhead camera. See Appendix D for details.
Figure 3: Autonomous rollouts of VITA across 5 AV-ALOHA tasks (CubeTransfer, SlotInsertion, HookPackage, PourTestTube, ThreadNeedle), and 2 Robomimic tasks (Square, Can). Notably, the AV-ALOHA tasks demand high-precision control, such as accurately pouring a small ball into a narrow tube opening, or threading a needle through a tiny hole.
@@
Figure 4: Autonomous rollouts of VITA on five challenging real-world tasks, including two bimanual AV-ALOHA tasks, HiddenPick, and TransferFromBox using active vision, and three single-arm ALOHA tasks, PickBall, ToothBrush and StoreDrawer
5.1 Experiment Settings
@@
Flow Matcher. We adopt OT-CFM (ot-cfm), which is based on optimal transport, and solve the ODE with an Euler solver using 6 linearly interpolated time steps $t$ for both VITA and FM.
Vision Encoding. We use ResNet-18 (resnet) as the vision encoder. A common practice in visuomotor policy learning is to use the vector-based visual features after global average pooling (su2025dense; chi2023diffusion), in which case VITA operates entirely on vector representations for both vision and action. We additionally evaluate a variant of VITA using the $9\times 512$ ResNet grid-based features to preserve more spatial information and assess VITA ’s scalability.
Baselines. Our baselines include state-of-the-art policies, including flow matching (FM) policy (fmp-affordance), diffusion policy (DP) (chi2023diffusion), and action chunking transformer (ACT) (zhao2023learning). We evaluate both efficiency and performance in Section˜5.2.
Training. We train VITA and baselines to predict action chunks of length 16, of which the first 8 actions are executed. We train VITA and FM on each task for 25K-50K steps. We follow the ACT and DP implementations in LeRobot (cadene2024lerobot). Since VITA and FM converge much faster than DP or ACT (a known advantage of FM methods (lipman2023flowmatchinggenerativemodeling)), we extend DP training to 100K steps, and ACT to 100K-200K steps. The training batch size is 128. See Appendix G for more detailed training settings.
Evaluation. We use 6 ODE steps for VITA and FM, and 10 DDPM (Denoising Diffusion Probabilistic Model) steps for DP (ho2020denoising). In simulation, we evaluate the policy every 500 training steps using 50 episodes per evaluation, and report the best success rate (SR) averaged over 3 seeds (Table˜2). For real-world tasks, we evaluate over 20 episodes per checkpoint on three single-arm tasks (Table˜3) and two bimanual manipulation tasks (Table˜4). Efficiency gains of VITA are analyzed in Section˜5.2.1, with all latency and memory measurements obtained on a single NVIDIA RTX 4090.
@@ 5.2 Performance
The main objective of this paper is to improve the time and space efficiency of visuomotor policies. This section demonstrates that VITA is a fast and precise visuomotor control policy, delivering efficiency gains (Section˜5.2.1) while matching or surpassing state-of-the-art policies (Section˜5.2.2).
5.2.1 Efficiency
FM and DP are typically parameterized using U-Nets (unet) or diffusion transformers (DiTs) (dit), which are trained to predict velocity fields or noise at each denoising step. U-Nets and DiTs are often large and computationally costly, which limits their real-time deployment that requires highly efficient inference. These architectures rely on explicit visual conditioning modules, e.g., cross-attention, FiLM, or AdaLN, which must be executed at every denoising step and inevitably increase both inference time and memory footprint.
A key determinant of efficiency in both conventional FM and VITA is the choice of visual latent representation. In conventional methods, vector-based and grid-based latents require different conditioning modules and thus incur different computational costs. In VITA, the visual and action latents must share the same dimensionality, so the latent choice directly determines the FM network architecture. We therefore analyze efficiency in these two settings separately. As shown below, VITA supports highly efficient architectures in both cases. We describe the implementations of VITA and conventional FM under each setting and explain why VITA yields efficiency advantages in both.
Vector-Based Visual Latents. With vector-based visual features, conventional methods typically employ FiLM (perez2018film) or AdaLN (dit) to condition the FM network. These methods compute modulation parameters via a separate conditioning network at every denoising step. FiLM modulates network outputs at each feature channel; AdaLN modulates normalization statistics at each network layer. The FM network is often implemented using transformers or U-Nets to effectively process noisy action chunks of shape $T_{\text{pred}}\times D_{\text{action}}$, and fuse in visual latents. In contrast, since VITA uses vector-based latents for both the vision source and action targets, the FM network $v_{\theta}$ reduces to a vector-to-vector mapping, with no need to fuse visual information. This enables a highly lightweight MLP-only architecture choice. We show that MLP-based VITA achieves better efficiency (both latency and memory) than the most efficient FM baseline (MLP-based), while matching the task performance of the strongest FM baseline (transformer-based).
Grid-Based Visual Latents. With grid-based features (e.g., $9\times 512$), cross-attention is often used in transformers to fuse visual tokens and action chunks for conditioning (dit-policy), which introduces quadratic time and space complexity. In contrast, VITA eliminates cross-attention in transformer-based implementations. As shown in Appendix B.6.2, VITA attains strong performance while being more efficient (Table˜1 grid-based settings).
To provide a comprehensive efficiency comparison, we implement two FM parameterizations (DiTs (dit) and U-Nets (unet)) and three conditioning mechanisms: AdaLN (dit) and FiLM (perez2018film) for vector-based features, and cross-attention (carp) for grid-based features.
We compare the inference latency, and inference memory usage of VITA and other FM based methods in Table˜1 for both vector-based and grid-based representations. VITA achieves an inference wall-clock time of 0.22 ms (vector-based) and 0.25 ms (grid-based) per action chunk, which is $1.5\times$ and $2\times$ faster than the best-performing FM baseline (transformer-based FM with similar model sizes) that incurs higher latency ($\sim$0.33 ms and $\sim$0.51 ms). VITA effectively reduces memory usage: the peak memory is 18.6% less than FM in the vector-based setting, and 28.7% less in the grid-based setting. Additionally, we discuss the training time and space efficiency in Appendix B.7. We further compare against FM accelerated with simple MLP architectures of similar model size to isolate the effect of architecture design in the vector-based setting. VITA remains $1.3\times$ faster at inference and uses 19.3% less memory, while FM with MLPs fails to achieve competitive success rates (Appendix B.6.1).
Beyond time and space efficiency, we show that VITA training enjoys faster convergence in Appendix Figure 15 compared to other policies.
Table 1: Comparison of the time and space efficiency of VITA and flow-matching baselines, grouped by the type of visual latents used (“Vector” or “Grid” based). Metrics include: model size, inference latency (ms/chunk, batch size 1), and inference memory (MiB), (see Appendix B.7.2 for inference memory measurement details).
Visual Model Architecture Conditioning Params Latency Memory
Vector VITA MLP N/A 31.09M 0.2215 333.86
FM Transformer AdaLN 31.16M 0.3307 410.38
FM U-Net FiLM 84.05M 0.3650 818.79
FM MLP AdaLN 32.20M 0.2831 413.95
DDPM U-Net FiLM 81.82M 2.5985 801.47
Grid VITA Transformer N/A 31.80M 0.2502 377.55
FM Transformer Cross-Attn 29.06M 0.5102 529.16
5.2.2 Success Rates
We evaluate VITA against state-of-the-art policies on 9 simulated tasks and 5 real-world tasks. We report SRs for FM using the transformer-based FM with AdaLN, as it achieves the strongest performance among the FM variants evaluated in Section˜5.2.1; other variants such as MLP-based FM perform poorly (Appendix B.6.1), while FM using cross-attention or U-Net reaches similar SRs but is substantially slower to train. Across the 9 simulation tasks (Table˜2) and the 3 real-world single-arm ALOHA tasks (Table˜3), VITA consistently outperforms or matches state-of-the-art methods. We further evaluate VITA on 2 challenging real-world bimanual AV-ALOHA tasks (Table˜4) featuring 21-DoF high-dimensional actions and active vision. We discuss the under-performance of DP in Appendix B.8.2, which is largely due to the stringent success criteria of multi-stage ALOHA tasks. Appendix E.1 demonstrates VITA’s robustness to online perturbations, and Appendix E.2 evaluates its generalization to unseen objects.
@@ 5.2.2 Success Rates
Table 2: SRs on simulation tasks on AV-ALOHA, Robomimic, PushT, and CloseBox. We report the mean and the standard deviation of the best SRs during validation across 3 random seeds.
@@
[…]
Table 3: Comparison of SRs on three real-world single-arm ALOHA manipulation tasks. Each task is decomposed into subtasks, and SRs are reported per subtask.
PickBall StoreDrawer ToothBrush
Pick Place Pick Place Close Pick Insert
VITA 0.75 0.70 1.00 0.95 0.95 0.80 0.50
FM 0.75 0.65 0.90 0.90 0.90 0.90 0.50
DP 0.60 0.60 1.00 0.95 0.90 0.60 0.30
ACT 0.50 0.45 0.65 0.65 0.50 0.70 0.30
Table 4: SRs of VITA on two real-world bimanual manipulation tasks with active vision on AV-ALOHA. Each task is decomposed into three subtasks, and SRs are reported per subtask.
@@
Reveal Pick Place Pick Transfer Place
VITA 1.00 0.65 0.65 1.00 0.95 0.90